OpenStudy (anonymous):

What is the opening direction of a parabola with the equation x=-y^2? PLEASE HELP

OpenStudy (campbell_st):

its concave left

OpenStudy (campbell_st):

and symmetrical about the y axis

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

notice that if the equation was x = y^2, x will ALWAYS be positve no matter what y is. so you equation is x = -y^2 means x will always be negative. so it opens left.

OpenStudy (anonymous):

can you guys solve this by graphing?? y = -x + 4 (x - 3)2 + y2 = 25

OpenStudy (anonymous):

y = -x + 4 (x - 3)2 + y2 = 25

OpenStudy (campbell_st):

yep... you need to solve this by substitution so take the 1st equation and substitute it into the 2nd equation \[(x - 3)^2 + (4 - x)^2 = 25\] 1st simplify the equation and the solve for x when you get x, there are 2 values...I think substitute into the 1st equation to find y

OpenStudy (anonymous):

it says to solve this system by graphing because I need the coordinates lol